OpenStudy (anonymous):

what is the midpoint of a segment with the endpoints a(5,-1), b(-3,7)

OpenStudy (anonymous):

{(X1+X2/2),(Y1+Y2)/2}

OpenStudy (anonymous):

\[\left\{ \frac{ {x_1}+{x_2} }{ 2 } , \frac{ {y_1}+{y_2} }{ 2 } \right\}\]

OpenStudy (anonymous):

what numbers do I put for x2,x1 and y1,y2? lol

OpenStudy (anonymous):

The first element is normally X. The order of -3 or 5 does not matter

OpenStudy (anonymous):

okay. what about for the y's? lol